# The PanAf-FGBG Dataset: Understanding the Impact of Backgrounds in Wildlife Behaviour Recognition

### v2

@@ Abstract
Computer vision analysis of camera trap video footage is essential for wildlife conservation, as captured behaviours offer some of the earliest indicators of changes in population health. Recently, several high-impact animal behaviour datasets and methods have been introduced to encourage their use; however, the role of behaviour-correlated background information and its significant effect on out-of-distribution generalisation remain unexplored. In response, we present the PanAf-FGBG dataset, featuring 20 hours of wild chimpanzee behaviours, recorded at over 350 individual camera locations. Uniquely, it pairs every video with a chimpanzee (referred to as a foreground video) with a corresponding background video (with no chimpanzee) from the same camera location. We present two views of the dataset: one with overlapping camera locations and one with disjoint locations. This setup enables, for the first time, direct evaluation of in-distribution and out-of-distribution conditions, and for the impact of backgrounds on behaviour recognition models to be quantified. All clips come with rich behavioural annotations and metadata including unique camera IDs and detailed textual scene descriptions. Additionally, we establish several baselines and present a highly effective latent-space normalisation technique that boosts out-of-distribution performance by +5.42% mAP for convolutional and +3.75% mAP for transformer-based models. Finally, we provide an in-depth analysis on the role of backgrounds in out-of-distribution behaviour recognition, including the so far unexplored impact of background durations (i.e., the count of background frames within foreground videos).
@@
Figure 1: Conceptual Overview. The PanAf-FGBG dataset comprises 20+ hours of paired and richly annotated foreground-background videos from camera traps in wild chimpanzee habitats (see example frames at the top). The dataset unlocks systematic analyses of the impact of background information on wildlife behaviour recognition. We provide baselines, quantify the background impact in relation to other sources of information, and demonstrate that utilising background information in latent space (see blue data point)) can significantly improve recognition performance in this challenging domain.
@@ 1 Introduction
In the human action recognition domain, poor OOD performance is well-known to originate from shortcut learning [14] of static cues or backgrounds [26, 20, 25]. This leads models to rely heavily on background information instead of the behaviour of interest, and several studies have shown that the majority of action recognition performance in this domain can be attributed to background learning [41, 18, 8]. Despite the known impact of backgrounds on action recognition and a critical need to advance model performance in challenging wildlife applications [40, 24], reliance on static cues for animal behaviour recognition has received only limited attention. This is largely due to a lack of datasets explicitly designed to study its effects.
@@
Our contributions are as follows: (i) we release PanAf-FGBG with its rich annotations and meta-data, providing a large wildlife dataset suitable for explicitly investigating the impact of backgrounds on behaviour recognition; (ii) we show models trained on background-only videos achieve strong animal behaviour recognition performance ($\sim$ 65% of baseline performance), corroborating findings in the human domain; (iii) we conduct the first evaluation of background duration in videos, revealing its impact on action recognition models and highlighting key differences between convolutional and attention-based networks; and (iv) we introduce a simple yet powerful latent-space background neutralisation that yields notable performance gains, particularly in OOD scenarios for both convolutional (+5.42% mAP) and transformer-based approaches (+3.75% mAP).
@@ 2 Related Work
Animal Behaviour Datasets. In recent years, several high-impact animal behaviour datasets have been introduced to the CV community [35, 6, 29, 31, 22, 4]. Although they all centre on animal behaviour understanding, they each have a different focus. The large-scale datasets Animal Kingdom [35] and MammalNet [6] focus on behavioural understanding across species, each including videos of $\sim$ 800 different species. They provide atomic action and high-level behaviour annotations, respectively, and support, in addition to action recognition, the related tasks of action localisation [35, 6] and pose estimation [35]. The smaller ChimpAct dataset [31] provides longitudinal data that follows a zoo-housed group of 20 chimpanzees with a focus on a single male juvenile. It also contains key-point pose information, and fine-grained spatio-temporal behaviour labels. While valuable, these datasets lack in-situ footage from ecological sensors like camera traps and drones, meaning animals are not observed in their natural environment. This may result in unrepresentative behaviour distributions and limit their suitability for studying wild behaviour [9, 5].
In contrast, KABR [22] and BaboonLand [13] comprise drone footage of free-living Kenyan wildlife, with fine-grained spatio-temporal behaviour annotations, focussing on individual and group-level behaviour, respectively. The LoTE dataset [29] provides camera trap video footage of 11 endangered species from a 12-year longitudinal study; it is accompanied by environmental information (i.e., weather conditions and habitat) and highlights the importance of temporal and spatial continuity for conservation efforts. The dataset most comparable to our own is the PanAf20k dataset [4], which comprises camera trap videos of great apes accompanied by multi-label behaviour annotations.
Our dataset orthogonally complements the growing body of work dedicated to understanding animal behaviour by providing the first dataset specifically designed to evaluate backgrounds and out-of-distribution generalisation for behaviour recognition. It is unique in providing foreground-background pairs, and its configurations for in- and out-of-distribution evaluation are defined identically to the well-known iWildCam  [3] and Wilds dataset [23] for species classification. Similar to LotE [29], we provide a detailed description of the background, however we also include a unique camera ID (i.e., geospatial location) to allow footage taken from the same camera to be linked. Similar to [4], our data is gathered from 14 national parks, spanning 6 different countries, although unlike ours they do not include metadata to allow videos to be linked to specific countries, research sites, locations, or habitats. Similarly, it is more diverse than [29] with its footage from a single national park.
@@
Background Bias Datasets. The effect of the background has often been modelled using synthetically generated data that removes either the foreground or the background. [15, 7, 8, 25]. For instance, Ilic et al. [20] present an synthetic version of the UCF101 dataset, animating spatial noise using image motion extracted by an optical flow estimator, Li et al. [25] replace the backgrounds of IID HMDB51 with sinusoidal stripe images, and Chung et al [8] adopt an in-painting model to remove actors from the video. Datasets that include fine-grained spatio-temporal action annotations also remove static cues implicitly by isolating the region of the video where the action is occurring.
@@ 3 Dataset
Dataset Overview and Statistics. PanAf-FGBG comprises footage gathered under ethical oversight as part of the The PanAf Programme: The Cultured Chimpanzee [1]. It contains 21 hours of camera trap footage of individual chimpanzees in tropical Africa. Its footage is collected from 389 individual camera locations across 14 national parks in 6 African countries. In total, we provide 5,070 video pairs, each 15-seconds in duration. All videos are accompanied by a multi-label behaviour annotation and scene metadata, plus the unique feature of foreground-background video pairing (see Fig. 1). Class and habitat distributions of the data are given in Fig. 2 and Fig. 3, respectively. Our dataset follows a long-tailed distribution common in animal behaviour datasets [35, 6, 29, 31, 4]. The three of the most commonly occurring classes are observed in $>$60% of videos, while the rarest are observed in fewer than $<$ 3%.
@@
Ethical Considerations. Both GPS location and habitat type for each camera was noted during this process. To protect the location of the apes from threats, such as poaching, we use randomly generated monikers for the names of the research sites and assign a unique camera ID to each video, instead of revealing the geospatial location of the cameras.
Data Annotation. The behavioural annotations were provided by users on the community science platform Chimp&See [2]. Annotators were presented with a choice of behaviour classification categories, chosen specifically by experts for their ecological importance. Behaviours not listed in the classification categories were also permitted. These annotations were then extracted and expertly grouped into 14 co-occurring classes, which form the multi-label behavioural annotations presented here. The annotations follow a multi-hot binary format that indicate the presence of one or more behaviours. It should also be noted that behaviours are not assigned to individual apes or temporally localised within each video. To ensure annotation quality and consistency a video was only deemed to be analysed when either three volunteers marked the video as blank, or unanimous agreement between seven volunteers was observed, or 15 volunteers annotated the video.
@@ 4 Preliminaries
We train multi-label classification models on both views of the dataset: $\mathcal{D}^{\text{overlap}}$ and $\mathcal{D}^{\text{disjoint}}$ using various combinations of foreground ($\mathcal{F}$), background ( $\mathcal{B}$) and synthetic background ($\tilde{\mathcal{B}}$) videos. The latter were created by generating a segmentation mask using SAM2 [37] followed by mean pixel value filling, similar to [19, 7, 8]. Both views are comprised of foreground-background video pairs $\{(v^{\mathcal{F}}_{i},v^{\mathcal{B}}_{i})\}_{i=1}^{N}$, where $v^{\mathcal{F}}_{i}\in\mathcal{F}$ and $v^{\mathcal{B}}_{i}\in\mathcal{B}$, with $N$ representing the total number of video pairs. Each foreground video $v^{\mathcal{F}}_{i}$ is associated with a video-level label vector $y_{i}=(y_{i1},y_{i2},\dots,y_{iC})$, where each $y_{ij}\in\{0,1\}$ indicates the presence ($1$) or absence (00) of the $j$-th class in video $i$, and $C$ is the total number of classes.
@@ 5.1 Setup
All experiments were performed using the ResNet-50 [17] and MViT-V2 [28] architectures, which were initialised with feature extractors pre-trained on Kinetics-400 [21]. For ResNet-50, both 2D and 3D convolution versions are tested. After initial experiments to ensure optimal optimiser-network pairings for the task, ResNet and MViT-V2 models were fine-tuned for 300 epochs using the SGD [38] and AdamW [30] optimisers, respectively. Both models followed a linear warm-up schedule followed by cosine annealing. For ResNet models, the learning rate was increased from $1\times 10^{-5}$ to $1\times 10^{-4}$ over 20 epochs, while it was increased from $1\times 10^{-6}$ to $1\times 10^{-4}$ over 30 epochs for MViT-V2 models. In both cases, momentum of 0.9 was utilised. All computation was performed on 4$\times$×NVIDIA H200 GPUs using a distributed batch size of 336. During training, video tensors with a spatio-temporal resolution of $16\times 256\times 256$ were employed. Frames were sampled uniformly at a rate of $1/24$, avoiding dense sampling strategies which are shown to exacerbate shortcut learning in action recognition models [8]. For MViT-V2 models, a patch size of $16\times 16$ pixels was utilised. Evaluation was performed using micro and mean Average Precision and are referred to as uAP and mAP, respectively.
@@ 5.2 Background Reliance
Experimental Overview. First, we examined the reliance of animal behaviour recognition on the background by training model architectures on the foreground ${\mathcal{F}}$, background ${\mathcal{B}}$, and synthetic background videos $\tilde{\mathcal{B}}$ only. We refer to models trained on ${\mathcal{F}}$ and ${\mathcal{B}}$ (or $\tilde{\mathcal{B}}$) as FG-only and BG-only models, respectively, as shown earlier in Fig. 1. To separate background information from class-distribution intrinsic information, a dummy classifier with no access to visual information was also evaluated. It would effectively return predictions based on the class distribution information alone. Since our dataset consists of unique foreground-background pairs, it was possible to train BG-only models by retaining the original labels associated with the foreground video of the pairing. In all cases, models were tested on ${\mathcal{F}}$ for comparability. Note that this makes the task more difficult for BG-only models, since animals are only observed at test time and not during training. All experiments are conducted on both $D^{overlap}$ as well as $D^{disjoint}$.
@@
Table 1: Quantifying Background Reliance. Performance comparison of models trained on $\mathcal{F}$, $\mathcal{B}$, and $\tilde{\mathcal{B}}$ videos. Dual-stream models are indicated by $\mathcal{F}+\mathcal{B}$. The absolute highest performance is indicated in bold whereas best performance using only a single-stream is underlined. The performance based on class distribution only is given by a dummy classifier (see row 1). The background-to-foreground performance ratio is also provided (see rows 6, 11, and 16). Performance is reported on $\mathcal{D}^{overlap}$ and $\mathcal{D}^{disjoint}$ where Fig. 1 visualises an mAP scatter plot between the two for 3D R50.
@@
Result 1 – Backgrounds are a Predictor of Animal Behaviour. Table 1 shows that BG-only models achieve strong performance when compared to their respective FG-only counterparts. To most clearly illustrate this, the ratio of background-to-foreground performance $\frac{AP^{B}}{AP^{F}}$ published in [8] is reported for each architecture in rows 6, 11, and 16. Background-to-foreground performance ratios are consistently high across all models, never dropping below 70% and 65% for uAP and mAP, respectively, on either dataset. This indicates that much of the behaviour recognition performance is indeed achievable by just utilisation of background information - the environment is a strong predictor of behaviour, even out-of-distribution. Thus, features of the environment are already behaviour predictors (trees for climbing, paths for travel, termite mound for feeding etc.). This result is further confounded by the fact that BG-only models outperform the non-visual classifier (row 1) by a significant margin, particularly on $\mathcal{D}^{overlap}$. On $\mathcal{D}^{disjoint}$ this observation still holds, although by somewhat smaller margins. Note that for the weakest model (2D R50) background mAP is even on-par or above the foreground baseline (see rows 2-3 in the final column). Together, these results illustrate powerfully, that across architectures background information alone is employable as a strong animal behaviour predictor.
@@
Result 2 – Transformers rely less on Backgrounds. On the overlapping dataset, the 3D-R50 and MViT-V2 models (rows 7-16) display significantly better performance when compared with the 2D-R50 (rows 2-6), and perform similarly to each other with respect to uAP and mAP. However, the MViT-V2 model widely outperforms the 3D R50 on the disjoint dataset, suggesting that it does not rely as heavily on background features for performance. In support of this, while all models experience a large decrease in performance when moving from the overlapping to the disjoint dataset (generalisation gap), the smallest impact is on the MViT-V2. These differences between architectures are highlighted most clearly by observing uAP on the disjoint dataset. Here, the background only ratio for the MViT-V2 model is $\sim$ 30% and $\sim$ 13% lower than the 2D-R50 and 3D-R50, respectively. MViT-V2 achieves the best overall performance and exhibits the lowest background-to-foreground performance ratio on both dataset configurations, corroborating similar findings from the still image domain that show transformers are more invariant to changes in the background [34].
@@
Analysing Background Information Contributions. The dual 3D-R50 model achieves significant performance gains over its FG-only baseline, particularly on $\mathcal{D}^{disjoint}$, where increases of 7.92% and 8.28% in uAP and mAP, respectively, can be observed. Smaller, but still significant improvements are also observed on $\mathcal{D}^{overlap}$ (+3.62% uAP and +4.25% mAP). Interestingly, the performance of the dual-stream 3D-R50 exceeds that of MViT-v2, including its dual-stream counterpart, on nearly all metrics except for disjoint uAP. The dual-stream MViT-V2 model, on the other hand, shows improved mAP but a lower uAP when compared with the baseline model, a trend consistent across both dataset configurations. This analysis indicates that BG-only performance in the wildlife domain is not a subset of the FG-only model – which of course also has access to background information. Instead, explicitly learned (and frozen) background features can aid the FG-only baseline in models such as ResNet via late fusion. Finally, all BG-only models trained on $\tilde{\mathcal{B}}$ seem to outperform those trained on $\mathcal{B}$. However, we will show later in Sec. 5.4 that non-synthetic background videos dominate synthetic counterparts when explicitly utilised for performance improvements above the foreground baselines.
Figure 5: The Effect of increasing Background Duration on Micro- and Mean- Average Precision. Comparison of 2D R50, 3D R50 and MViT-V2 model performance as a function of the simulated background duration ($\lambda$). The 3D-R50 is more sensitive to increasing background durations than MViT-V2. MViT-V2 is relatively robust to increasing background durations on the overlapping dataset, although it becomes highly sensitive when evaluated on OOD data.

### v3

@@ Abstract
Computer vision analysis of camera trap video footage is essential for wildlife conservation, as captured behaviours offer some of the earliest indicators of changes in population health. Recently, several high-impact animal behaviour datasets and methods have been introduced to encourage their use; however, the role of behaviour-correlated background information and its significant effect on out-of-distribution generalisation remain unexplored. In response, we present the PanAf-FGBG dataset, featuring 21 hours of wild chimpanzee behaviours, recorded at 389 individual camera locations. Uniquely, it pairs every video with a chimpanzee (referred to as a foreground video) with a corresponding background video (with no chimpanzee) from the same camera location. We present two views of the dataset: one with overlapping camera locations and one with disjoint locations. This setup enables, for the first time, direct evaluation of in-distribution and out-of-distribution conditions, and for the impact of backgrounds on behaviour recognition models to be quantified. All clips have rich behavioural annotations and metadata, including unique camera IDs. Additionally, we establish several baselines and present a highly effective latent-space normalisation technique that boosts out-of-distribution performance by +5.42% mAP for convolutional and +3.75% mAP for transformer-based models. Finally, we provide an in-depth analysis on the role of backgrounds in out-of-distribution behaviour recognition, including the so far unexplored impact of background durations. (i.e., the count of background frames within foreground videos). The dataset is available at https://obrookes.github.io/panaf-fgbg.github.io/.
@@
Figure 1: Conceptual Overview. The PanAf-FGBG dataset comprises $>$20 hours of paired and richly annotated foreground-background camera trap videos of wild chimpanzees. The dataset unlocks systematic analyses of the impact of background information on wildlife behaviour recognition. We provide baselines, quantify the background impact on performance, and demonstrate that utilising background information in latent space (see blue data point) can significantly improve recognition performance in this challenging domain.
@@ 1 Introduction
In the human action recognition domain, poor OOD performance is well-known to originate from shortcut learning [18] of static cues or backgrounds [30, 24, 29]. This leads models to rely heavily on background information instead of the behaviour of interest solely [2], and several studies have shown that the majority of action recognition performance in this domain can be attributed to background learning [49, 22, 12]. Despite the known impact of backgrounds on action recognition and a critical need to advance model performance in wildlife applications [48, 28, 42], reliance on static cues for animal behaviour recognition has received limited attention. This is largely due to a lack of datasets explicitly designed to study its effects.
@@
Our contributions are as follows: (i) we release PanAf-FGBG with its rich annotations and meta-data, providing a large wildlife dataset suitable for explicitly investigating the impact of backgrounds on behaviour recognition; (ii) we show models trained on background-only videos achieve strong animal behaviour recognition performance ($\sim$ 65% of baseline performance), corroborating findings in the human domain; (iii) we conduct the first evaluation of background duration in videos, revealing its impact on action recognition models and highlighting key differences between convolutional and attention-based networks; and (iv) we introduce a simple yet powerful latent-space background neutralisation that yields notable performance gains, particularly in OOD scenarios for convolutional (+5.42% mAP) and transformer-based approaches (+3.75% mAP).
@@ 2 Related Work
Animal Behaviour Datasets. In recent years, several high-impact animal behaviour datasets have been introduced to the CV community, often accompanied by new methods [46, 4, 40, 10, 33, 6, 26, 8, 7]. Although they all centre on animal behaviour understanding, they each have a different focus. The large-scale datasets Animal Kingdom [40] and MammalNet [10] concentrate on behavioural understanding across species, each including videos of $\sim$ 800 different species. They provide atomic action and high-level behaviour annotations, respectively, and support, in addition to action recognition, the related tasks of action localisation [40, 10] and pose estimation [40]. The smaller ChimpAct dataset [36] provides longitudinal data that follows a zoo-housed group of 20 chimpanzees with a focus on a single male juvenile. It also contains key-point pose information, and fine-grained spatio-temporal behaviour labels. While valuable, these datasets lack in-situ footage from ecological sensors like camera traps and drones, meaning animals are not observed in their natural environment. This may result in unrepresentative behaviour distributions and limit their suitability for studying wild behaviour [13, 9].
In contrast, KABR [26] and BaboonLand [17] comprise drone footage of free-living Kenyan wildlife, with fine-grained spatio-temporal behaviour annotations, focussing on individual and group-level behaviour, respectively. The LoTE dataset [33] provides camera trap video footage of 11 endangered species from a 12-year longitudinal study; it is accompanied by environmental information (i.e., weather conditions and habitat) and highlights the importance of temporal and spatial continuity for conservation efforts. The dataset most comparable to our own is the PanAf20k dataset [8] and its related PanAf500-based works [52, 46, 53]. PanAf20k comprises camera trap videos of great apes accompanied by multi-label behaviour annotations.
Our dataset orthogonally complements this growing body of work dedicated to understanding animal behaviour by providing the first dataset specifically designed to evaluate backgrounds and out-of-distribution generalisation for behaviour recognition. It is unique in providing foreground-background pairs, and its configurations for in- and out-of-distribution evaluation are defined identically to the well-known iWildCam  [5] and Wilds dataset [27] for species classification. Similar to LotE [33], we provide a detailed description of the background, however we also include a unique camera ID (i.e., geospatial location) to allow footage taken from the same camera to be linked. Similar to [8], our data is gathered from 14 national parks, spanning 6 different countries, although unlike ours they do not include metadata to allow videos to be linked to specific countries, research sites, locations, or habitats. Similarly, it is more diverse than [33] with its footage from a single national park.
@@
Background Bias Datasets. The effect of the background has often been modelled using synthetically generated data that removes either the foreground or the background. [19, 11, 12, 29]. For instance, Ilic et al. [24] present a synthetic version of the UCF101 dataset, animating spatial noise using image motion extracted by an optical flow estimator, Li et al. [29] replace the backgrounds of IID HMDB51 with sinusoidal stripe images, and Chung et al [12] adopt an in-painting model to remove actors from the video. Datasets that include fine-grained spatio-temporal action annotations also remove static cues implicitly by isolating the region of the video where the action is occurring.
@@ 3 Dataset
Dataset Overview and Statistics. PanAf-FGBG comprises footage gathered under ethical oversight as part of the PanAf Programme: The Cultured Chimpanzee [1]. It contains 21 hours of camera trap footage of individual chimpanzees in tropical Africa. Its footage is collected from 389 individual camera locations across 14 national parks in 6 African countries. In total, we provide 5,070 video pairs, each 15 seconds in duration. All videos are accompanied by a multi-label behaviour annotation and scene metadata, plus the unique feature of foreground-background video pairing (see Fig. 1). Class and habitat distributions of the data are given in Fig. 2 and Fig. 3, respectively. Our dataset follows a long-tailed distribution [34, 41] common in animal behaviour datasets [40, 10, 33, 36, 8]. The three of the most commonly occurring classes are observed in $>$60% of videos, while the rarest are observed in fewer than $<$ 3%.
@@
Ethical Considerations. To protect the location of the apes from threats (e.g., poaching) we use monikers for the names of the research sites and assign a unique camera ID to each video, hiding the geospatial location of the cameras.
Data Annotation. The behavioural annotations were provided by users on the community science platform Chimp&See [3]. Annotators were presented with a choice of behaviour classification categories, chosen specifically by experts for their ecological importance. Behaviours not listed in the classification categories were also permitted. These annotations were then extracted and expertly grouped into 14 co-occurring classes, which form the multi-label behavioural annotations presented here. The annotations follow a multi-hot binary format that indicate the presence of one or more behaviours. It should also be noted that behaviours are not assigned to individual apes or temporally localised within each video. To ensure annotation quality and consistency a video was only deemed to be analysed when either three volunteers marked the video as blank, or unanimous agreement between seven volunteers was observed, or 15 volunteers annotated the video. Habitat type was noted during camera setup.
Foreground-Background Pair Selection. Pairs are formed by selecting the temporally closest available foreground and background videos to ensure similarity. Complete standardisation statistics can be found in Sup. Mat.
@@ 4 Preliminaries
We train multi-label classification models on both views of the dataset: $\mathcal{D}^{\text{overlap}}$ and $\mathcal{D}^{\text{disjoint}}$ using various combinations of foreground ($\mathcal{F}$), background ( $\mathcal{B}$) and synthetic background ($\tilde{\mathcal{B}}$) videos. The latter were created by generating a segmentation mask using SAM2 [44] followed by mean pixel value filling, similar to [23, 11, 12]. Both views comprise foreground-background video pairs $\{(v^{\mathcal{F}}_{i},v^{\mathcal{B}}_{i})\}_{i=1}^{N}$, where $v^{\mathcal{F}}_{i}\in\mathcal{F}$ and $v^{\mathcal{B}}_{i}\in\mathcal{B}$, with $N$ representing the total number of video pairs. Each foreground video $v^{\mathcal{F}}_{i}$ is associated with a video-level label vector $y_{i}=(y_{i1},y_{i2},\dots,y_{iC})$, where each $y_{ij}\in\{0,1\}$ indicates the presence ($1$) or absence (00) of the $j$-th class in video $i$, and $C$ is the total number of classes.
@@ 5.1 Setup
All experiments were performed using the ResNet-50 [21] and MViT-V2 [32] architectures, which were initialised with feature extractors pre-trained on Kinetics-400 [25]. For ResNet-50, both 2D and 3D convolution versions are tested. After initial experiments to ensure optimal optimiser-network pairings for the task, ResNet and MViT-V2 models were fine-tuned for 300 epochs using the SGD [45] and AdamW [35] optimisers, respectively. Both models followed a linear warm-up schedule with subsequent cosine annealing. For ResNet models, the learning rate was increased from $1\times 10^{-5}$ to $1\times 10^{-4}$ over 20 epochs, while it was increased from $1\times 10^{-6}$ to $1\times 10^{-4}$ over 30 epochs for MViT-V2 models. In both cases, momentum of 0.9 was utilised. All computation was performed on 4$\times$×NVIDIA H200 GPUs using a distributed batch size of 336. During training, video tensors with a spatio-temporal resolution of $16\times 256\times 256$ were employed. Frames were sampled uniformly at a rate of $1/24$, avoiding dense sampling strategies which are shown to exacerbate shortcut learning in action recognition models [12]. For MViT-V2 models, a patch size of $16\times 16$ pixels was utilised. Evaluation was performed using micro and mean Average Precision and are referred to as uAP and mAP, respectively.
@@ 5.2 Background Reliance
Experimental Overview. First, we examined the reliance of animal behaviour recognition on the background by training model architectures on the foreground ${\mathcal{F}}$, background ${\mathcal{B}}$, and synthetic background videos $\tilde{\mathcal{B}}$ only. We refer to models trained on ${\mathcal{F}}$ and ${\mathcal{B}}$ (or $\tilde{\mathcal{B}}$) as FG-only and BG-only models, respectively, as shown earlier in Fig. 1. To separate background information from class-distribution intrinsic information, a dummy classifier with no access to visual information was also evaluated. It would effectively return predictions based on the class distribution information alone. Since our dataset consists of unique foreground-background pairs, it was possible to train BG-only models by retaining the original labels associated with the foreground video of the pairing. In all cases, models were tested on ${\mathcal{F}}$ for comparability. Note that this makes the task more difficult for BG-only models, since animals are only observed at test time and not during training. All experiments are conducted on both $D^{overlap}$ and $D^{disjoint}$.
@@
Table 1: Quantifying Background Reliance. Performance comparison of models trained on $\mathcal{F}$, $\mathcal{B}$, and $\tilde{\mathcal{B}}$ videos. Dual-stream models are indicated by $\mathcal{F}+\mathcal{B}$. The absolute highest performance is indicated in bold whereas best performance using only a single-stream is underlined. The performance based on class distribution only is given by a dummy classifier (see row 1). The background-to-foreground performance ratio ($\frac{AP^{B}}{AP^{F}}$) is also provided (see rows 6, 11, and 16). Performance is reported on $\mathcal{D}^{overlap}$ and $\mathcal{D}^{disjoint}$ where Fig. 1 visualises an mAP scatter plot between the two for 3D R50.
@@
Result 1 – Backgrounds are a Predictor of Animal Behaviour. Table 1 shows that BG-only models achieve strong performance when compared to their respective FG-only counterparts. To illustrate this, the ratio of background-to-foreground performance $\frac{AP^{B}}{AP^{F}}$ published in [12], showing how close a model can get to FG performance with just BG cues, is reported for each architecture in rows 6, 11, and 16. Background-to-foreground performance ratios are consistently high across all models, never dropping below 0.70 and 0.65 for uAP and mAP, respectively, on either dataset. This indicates that much of the behaviour recognition performance is indeed achievable by just utilisation of background information - the environment is a strong predictor of behaviour, even out-of-distribution. Thus, features of the environment are already behaviour predictors (trees for climbing, paths for travel, termite mound for feeding etc.). This result is further compounded by the fact that BG-only models outperform the non-visual classifier (row 1) by a significant margin, particularly on $\mathcal{D}^{overlap}$. On $\mathcal{D}^{disjoint}$ this observation still holds, although by somewhat smaller margins. Note that for the weakest model (2D R50) background mAP is even on-par or above the foreground baseline (see rows 2-3 in the final column). Together, these results illustrate powerfully, that across architectures background information alone is employable as a strong animal behaviour predictor.
@@
Result 2 – Multi-scale Vision Transformers rely less on backgrounds. On the overlapping dataset, the 3D-R50 and MViT-V2 models (rows 7-16) display significantly better performance when compared with the 2D-R50 (rows 2-6), and perform similarly to each other with respect to uAP and mAP. However, the MViT-V2 model widely outperforms the 3D R50 on the disjoint dataset, suggesting that it does not rely as heavily on background features for performance. In support of this, while all models experience a large decrease in performance when moving from the overlapping to the disjoint dataset (generalisation gap), the smallest impact is on the MViT-V2. These differences between architectures are highlighted most clearly by observing uAP on the disjoint dataset. Here, the background only ratio for the MViT-V2 model is $\sim$ 30% and $\sim$ 13% lower than the 2D-R50 and 3D-R50, respectively. MViT-V2 achieves the best overall performance and exhibits the lowest background-to-foreground performance ratio on both dataset configurations, corroborating similar findings from the still image domain that show transformers are more invariant to changes in the background [39].
@@
Result 3 - Analysing Background Information Contributions. The dual 3D-R50 model achieves significant performance gains over its FG-only baseline, particularly on $\mathcal{D}^{disjoint}$, where increases of 7.92% and 8.28% in uAP and mAP, respectively, can be observed. Smaller, but still significant improvements are also observed on $\mathcal{D}^{overlap}$ (+3.62% uAP and +4.25% mAP). Interestingly, the performance of the dual-stream 3D-R50 exceeds that of MViT-v2, including its dual-stream counterpart, on nearly all metrics except for disjoint uAP. The dual-stream MViT-V2 model, on the other hand, shows improved mAP but a lower uAP when compared with the baseline model, a trend consistent across both dataset configurations. This analysis indicates that BG-only performance in the wildlife domain is not a subset of the FG-only model – which of course also has access to background information. Instead, explicitly learned (and frozen) background features can aid the FG-only baseline in models such as ResNet via late fusion. Finally, all BG-only models trained on $\tilde{\mathcal{B}}$ seem to outperform those trained on $\mathcal{B}$. However, we will show later in Sec. 5.4 that non-synthetic background videos dominate synthetic counterparts when explicitly utilised for performance improvements above the foreground baselines.